OpenStudy (anonymous):

@ivettef365 how would I find the slope intercept form, and point slope form of: Passes through (-1, -1) and is perpendicular to y = x - 8.

OpenStudy (ivettef365):

ok to be perpendicular the slope needs to be opposite reciprocal so you have y = x - 8 , slope = 1, so opposite reciprocal would be -1 right??

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so the slope has to equal -1

OpenStudy (ivettef365):

now use the formula to find b, so you have y = mx + b where y=-1. x=-1 and m = -1, so what is b

OpenStudy (anonymous):

-2 @ivettef365

OpenStudy (anonymous):

so my slope intercept formula would be y=-1*-1+-2?

OpenStudy (anonymous):

no y=-1x-2?

OpenStudy (ivettef365):

right that is correct, and you can write it as follows: y = -x -2

OpenStudy (anonymous):

ok and what about the point slope form, y+1=-1(x+1)

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!